# Giant bubbles of Fisher zeros in the quantum XY chain

### v1

@@ Abstract
The idea of complex $\beta$ goes back to Fisher who, following the pioneering work of Lee and Yang [15, 16], showed that the location of zeros of $Z$ on the complex $\beta$ plane, as the solution of $Z(\beta=\beta_{r}+i\beta_{i})=0$ and now referred to as Fisher zeros, can be used to identify thermodynamic phase transitions [17]. Later, in the study of quench dynamics from a given pure state, the zeros of the so-called boundary partition function (analogous to $Z$) were used to describe dynamical quantum phase transitions [18]. We have shown that the complex $Z$ and its Fisher zeros can be applied to quantum many-body systems to characterize quantum criticality [19, 20], low-energy excitations [21], breakdowns of the eigenstate thermalization hypothesis [22], and even topological features [23]. By now, there is a surge of interest in complexified quantum many-body systems [24, 21, 25, 26, 19, 20, 27, 28, 22, 23, 29, 30, 31]. In this context, the best understood example of Fisher zeros is the the one-dimensional transverse field Ising model (TFIM). For example, an open line of Fisher zeros extending to $\beta_{r}\rightarrow\infty$ in the complex $\beta$-plane indicates low energy excitations, such as domain-walls, at energy scale set by $1/\beta_{i}$. A natural question then is what happens to the configuration of Fisher zeros if a system does not have well-defined quasiparticles but an enhanced, broad, incoherent spectral weight (the suppressed spectral weight then qualifies as a pseudogap)? Clearly, they should not form open lines. To address these questions, we need to go beyond TFIM to study models that feature interesting, unconventional phases that do not resemble a Fermi liquid.
@@
Here $\sigma^{\mu}_{j}~(\mu=x,y,z)$ are Pauli matrices at site $j$, $g$ is the transverse field, $\gamma$ is the anisotropy parameter, and periodic boundary condition is assumed. As a paradigmatic example of quantum magnetism, this model has a rich phase diagram [38, 39, 40] summarized in Fig. 1. The $g=1$ line separates the disordered ($g>1$) from the ordered ferromagnetic phase ($g<1$). But the region bounded by the line $\gamma^{2}+g^{2}=1$ is peculiar: its spin correlation function oscillates on top of power law decay, and for finite system sizes, the gap also exhibits oscillations. As illustrated in Fig. 1 for $L=16$, the gap vanishes along a set of lines in the $g-\gamma$ plane, which become increasingly dense with larger system sizes [40]. This region is usually referred to as the oscillatory or incommensurate phase. The model reduces to TFIM along the line $\gamma=1$. And for $\gamma=0$, it reduces to the $XX$ model which serves as a representative example of a gapless Luttinger liquid in the thermodynamic limit.

### v2

@@ Abstract
The idea of complex $\beta$ goes back to Fisher who, following the pioneering work of Lee and Yang [15, 16], showed that the location of zeros of $Z$ on the complex $\beta$ plane, as the solution of $Z(\beta=\beta_{r}+i\beta_{i})=0$ and now referred to as Fisher zeros, can be used to identify thermodynamic phase transitions [17]. Later, in the study of quench dynamics from a given pure state, the zeros of the so-called boundary partition function (analogous to $Z$) were used to describe dynamical quantum phase transitions [18]. We have shown that the complex $Z$ and its Fisher zeros can be applied to quantum many-body systems to characterize quantum criticality [19, 20], low-energy excitations [21], breakdowns of the eigenstate thermalization hypothesis [22], and even topological features [23]. By now, there is a surge of interest in complexified quantum many-body systems [24, 21, 25, 26, 19, 20, 27, 28, 29, 22, 23, 30, 31, 32]. In this context, the best understood example of Fisher zeros is the the one-dimensional transverse field Ising model (TFIM). For example, an open line of Fisher zeros extending to $\beta_{r}\rightarrow\infty$ in the complex $\beta$-plane indicates low energy excitations, such as domain-walls, at energy scale set by $1/\beta_{i}$. A natural question then is what happens to the configuration of Fisher zeros if a system does not have well-defined quasiparticles but an enhanced, broad, incoherent spectral weight (the suppressed spectral weight then qualifies as a pseudogap)? Clearly, they should not form open lines. To address these questions, we need to go beyond TFIM to study models that feature interesting, unconventional phases that do not resemble a Fermi liquid.
@@
Here $\sigma^{\mu}_{j}~(\mu=x,y,z)$ are Pauli matrices at site $j$, $g$ is the transverse field, $\gamma$ is the anisotropy parameter, and period boundary condition is assumed. As a paradigmatic example of quantum magnetism, this model has a rich phase diagram [39, 40, 41] summarized in Fig. 1. The $g=1$ line separates the disordered ($g>1$) from the ordered ferromagnetic phase ($g<1$). But the region bounded by the line $\gamma^{2}+g^{2}=1$ is peculiar: its spin correlation function oscillates on top of power law decay, and for finite system sizes, the gap also exhibit oscillations. As illustrated in Fig. 1 for $L=16$, the gap vanishes along a set of lines in the $g-\gamma$ plane, which become increasingly dense with larger system sizes [41]. This region is usually referred to as the oscillatory or incommensurate phase. The model reduces to TFIM along the line $\gamma=1$. And for $\gamma=0$, it reduces to the $XX$ model which serves as a representative example of a gapless Luttinger liquid in the thermodynamic limit.